OpenStudy (anonymous):

Solve the equation. -9 - b + 8b = -23

OpenStudy (anonymous):

First step, group the like terms

OpenStudy (anonymous):

-b + 8b becomes 7b

OpenStudy (anonymous):

Then add 9 to each side. This will get the unknown by itself on the left side

OpenStudy (anonymous):

From here you can divide each side by the coefficient to get the unknown , b, all by itself and equal to your answer

OpenStudy (anonymous):

-9-b+8b=-23 -9+7b=-23 7b=-23 +9 7b=-14 b=-2